# IFT-UAM/CSIC-25-76

### v1

@@ Abstract
evortran is a modern Fortran library designed for high-performance genetic algorithms and evolutionary optimization. evortran can be used to tackle a wide range of problems in high-energy physics and beyond, such as derivative-free parameter optimization, complex search taks, parameter scans and fitting experimental data under the presence of instrumental noise. The library is built as an fpm package with flexibility and efficiency in mind, while also offering a simple installation process, user interface and integration into existing Fortran programs. evortran offers a variety of selection, crossover, mutation and elitism strategies, with which users can tailor an evolutionary algorithm to their specific needs. evortran supports different abstraction levels: from operating directly on individuals and populations, to running full evolutionary cycles, and even enabling migration between independently evolving populations to enhance convergence and maintain diversity. In this paper, we present the functionality of the evortran library, demonstrate its capabilities with example benchmark applications, and compare its performance with existing genetic algorithm frameworks. As physics-motivated applications, we use evortran to confront extended Higgs sectors with LHC data and to reconstruct gravitational wave spectra and the underlying physical parameters from LISA mock data, demonstrating its effectiveness in realistic, data-driven scenarios.
@@ 1 Introduction
While several GA libraries with varying degree of flexibility exist, evortran was developed to provide a modern Fortran-based alternative that balances performance and ease of use. Like Pikaia, it is available as a Fortran package manager (fpm[51, 52] package, making the compilation, installation and integration into other Fortran programs seamless. Compared to C/C++ libraries, evortran offers a more user friendly interface, while still making use of the increased computational performance of a low-level programming language, outperforming native Python implementations. One of the key strengths of GAs is their inherently modular structure, which allows for straightforward and efficient parallelization, operating on individuals in parallel across one or more populations. To take advantage of this property, evortran includes support for parallel execution using OpenMP, which significantly reduces the runtime on multi-core systems. This makes evortran and attractive choice for computationally intensive optimization problems in scientific computing. Moreover, evortran is designed with a modular structure that allows new GA operations for selection, crossover, mutation and selection to be easily added, such that the library can naturally grow and evolve over time. Finally, evortran stands out due to its flexible design which allows users to work at different abstraction levels: (1) the user can operate on individuals directly, (2) the user can evolve whole populations, (3) and the user can evolve a set of populations in parallel with the possibility of periodically migrating individuals in between the populations.
@@ 2.3.1 Selection routines
roulette-wheel I/F wheele_size [3]
@@ 2.3.2 Crossover routines
The blend crossover routine (also called BLX-$\alpha$) is a crossing procedure that can only be applied to individuals with genes consisting of continuous numbers and not integers. This method generates two offspring individuals by selecting new gene values within an extended range between the gene values of two parent individuals. With $a_{i}$ being the genes from parent A and $b_{i}$ the genes from parent B, the genes $c_{i}$ and $d_{i}$ of the offspring individuals C and D are given, respectively, by randomly and uniformly selecting a number in the ranges
@@
$c_{i}$ $=0.5\left[(1+\beta_{q,i})a_{i}+(1-\beta_{q,i})b_{i}\right]\,,$ (6)
$d_{i}$ $=0.5\left[(1-\beta_{q,i})a_{i}+(1+\beta_{q,i})b_{i}\right]\,.$
@@ 2.4 Evolution of a population
  • elitism: Elitism mode, where currently only the default option best is supported, see Section 2.3.4.
@@
  • offspring_size: Number of offspring individuals to generate.
@@
  • mutate: The mutation method, currently possible values are uniform and shuffle for both integer and float populations, and additionally gaussian only for float populations, see Section 2.3.3. (default is uniform.
@@ 3.2 Quick start: basic usage and main features
  • The helper subroutine get_xy rescales the genes values from the normalized interval $[0,1]$ to the domain $-2\leq x\leq 2$ and $-1\leq y\leq 3$ in which the global minimum should be determined by the algorithm.
@@ 4.1.2 Michalewicz function
{adjustwidth}
1cm
1 best_ind = evolve_population( &
@@ 4.1.3 Himmelblau function
$(x,y)$ (3.0, 2.0),absent3.02.0$\approx$(3.0,\ 2.0),≈ ( 3.0 , 2.0 ) , (11)
$(x,y)$ (2.805, 3.131),absent2.8053.131$\approx$(-2.805,\ 3.131),≈ ( - 2.805 , 3.131 ) , (12)
$(x,y)$ $\approx(-3.779,\ -3.283),$ (13)
$(x,y)$ $\approx(3.584,\ -1.848),$ (14)
@@ 4.1.4 Drop-Wave function
To demonstrate the flexibility of the evortran library, and to compare the performance of different crossover methods, we conducted a comprehensive set of runs using the evolve_migration function to minimze the Drop-Wave function using different crossover routines. For each dimensionality from $n=2$ to $n=10$, we tested each crossover operator (blend, sbx, one-point, two-point, uniform) in combination with all possible permutations of the three available selection routines (tournament, rank, roulette) and the three mutation strategies (uniform, shuffle, gaussian). Each unique combination was run 10 times, resulting in a total of $3\cdot 3\cdot 10=90$ runs per crossover method and dimension. The GA was configured via the following call:
{adjustwidth}
1cm
@@ 4.2.1 Confronting extended Higgs sector with LHC data
    Perturbative unitarity: We verify that the leading-order $2\to 2$ scalar scattering amplitudes remain below the threshold $8\pi$ in the high-energy limit. This results in a set of inequalities on combinations of the scalar quartic couplings, whose precise form can be found in Ref. [16].
@@
Since the scan covers all physically distinct combinations of the three scalar mixing angles $\alpha_{1,2,3}$, see Table 4, no distinction between $h_{2}$ and $h_{3}$ is imposed on their possible masses and couplings. As a result, the orange points corresponding to the state $h_{2}$ and the green points corresponding to the state $h_{3}$ show a similar distribution in the two plots of Fig. 8. Furthermore, the mass ranges for $h_{2}$ and $h_{3}$ span from about 63 GeV to 1 TeV. The lower bound arises because masses below $62.5$ GeV would open the decay channels $h_{1}\to h_{2}h_{2}/h_{3}h_{3}$, which tends to spoil the Standard-Model-like nature of $h_{1}$ unless the corresponding couplings governing these decays are fine-tuned to suppress the decays. The GA does not converge to such tuned regions of parameter space in this broad scan. On the upper end, the scan reaches the maximum mass range of 1 TeV set in the input.
@@
In the right plot of Fig. 9, we show the values of the coupling modifiers $\kappa_{u}$ vs. $\kappa_{d}$, which specifically correspond to the ratios of the $h_{1}$ couplings to top and bottom quarks, respectively, compared to the Standard Model. Unlike in the Yukawa Type I, these two couplings are now controlled independently via the alignment parameters $\xi_{u}$ and $\xi_{d}$. The distribution shows that most viable parameter points cluster in the region where both $\kappa_{u}$ and $\kappa_{d}$ are close to either $+1$ or $-1$, reflecting Standard-Model-like behavior. However, one point clearly stands out as it corresponds to a parameter point where $\kappa_{u}$ and $\kappa_{d}$ have opposite signs. The two points in the left and the right plot of Fig. 9 corresponding to this parameter point are highlighted in orange. This sign configuration between $\kappa_{u}$ and $\kappa_{d}$ (also referred to as wrong-sign Yukawa coupling regime) is a distinctive feature of the flavor-aligned S2HDM compared to the S2HDM Type I and can, as observed here, still be consistent with current data, though it is associated with modified interference effects in loop-induced Higgs processes (such as gluon fusion production or the $h\to\gamma\gamma$ decay).
@@ 5 Conclusions
With its modular design and simple user interface, evortran enables users to easily apply evolutionary strategies to complex optimization problems, including those with non-differentiable, discontinuous, or noisy fitness functions. The library supports customization of GA components, native real and integer encodings, and parallel execution via OpenMP. evortran is installed with the Fortran package manager fpm, ensuring a straightforward dependency management, compilation and installation process, as well as a seamless integration into both simple scripts and larger code bases.

### v2

@@ Abstract
evortran is a modern Fortran library designed for high-performance genetic algorithms and evolutionary optimization. evortran can be used to tackle a wide range of problems in high-energy physics and beyond, such as derivative-free parameter optimization, complex search taks, parameter scans and fitting experimental data under the presence of instrumental noise. The library is built as an fpm package with flexibility and efficiency in mind, while also offering a simple installation process, user interface and integration into existing Fortran (or Python) programs. evortran offers a variety of selection, crossover, mutation and elitism strategies, with which users can tailor an evolutionary algorithm to their specific needs. evortran supports different abstraction levels: from operating directly on individuals and populations, to running full evolutionary cycles, and even enabling migration between independently evolving populations to enhance convergence and maintain diversity. In this paper, we present the functionality of the evortran library, demonstrate its capabilities with example benchmark applications, and compare its performance with existing genetic algorithm frameworks. As physics-motivated applications, we use evortran to confront extended Higgs sectors with LHC data and to reconstruct gravitational wave spectra and the underlying physical parameters from LISA mock data, demonstrating its effectiveness in realistic, data-driven scenarios.
@@ 1 Introduction
While several GA libraries with varying degree of flexibility exist, evortran was developed to provide a modern Fortran-based alternative that balances performance and ease of use. Like Pikaia, it is available as a Fortran package manager (fpm[51, 52] package, making the compilation, installation and integration into other Fortran programs seamless. Compared to C/C++ libraries, evortran offers a more user-friendly interface, while still making use of the increased computational performance of a low-level programming language. In addition, its main optimization routines can be accessed from Python via lightweight wrappers, allowing seamless integration into Python-based workflows and outperforming native Python implementations. One of the key strengths of GAs is their inherently modular structure, which allows for straightforward and efficient parallelization, operating on individuals in parallel across one or more populations. To take advantage of this property, evortran includes support for parallel execution using OpenMP, which significantly reduces the runtime on multi-core systems. This makes evortran and attractive choice for computationally intensive optimization problems in scientific computing. Moreover, evortran is designed with a modular structure that allows new GA operations for selection, crossover, mutation and selection to be easily added, such that the library can naturally grow and evolve over time. Finally, evortran stands out due to its flexible design which allows users to work at different abstraction levels: (1) the user can operate on individuals directly, (2) the user can evolve whole populations, (3) and the user can evolve a set of populations in parallel with the possibility of periodically migrating individuals in between the populations.
@@ 2.3.1 Selection routines
roulette I/F wheele_size [3]
@@ 2.3.2 Crossover routines
[…]
@@
[…]
@@ 2.4 Evolution of a population
  • elitism: Elitism mode, where currently only the default option best_fitness is supported, see Section 2.3.4.
@@
  • offspring_size: Number of offspring individuals to generate (default is pop_size).
@@
  • mutate: The mutation method, currently possible values are uniform and shuffle for both integer and float populations, and additionally gaussian only for float populations, see Section 2.3.3. (default is uniform).
@@ 3.2 Quick start: basic usage and main features
  • The helper subroutine get_xy rescales the genes values from the normalized interval [0,1][0,1][ 0 , 1 ] to the domain $-2\leq x\leq 2$ and $-1\leq y\leq 3$ in which the global minimum should be determined by the algorithm.
@@ 1 Rosenbrock function:
3.3 Using evortran from Python
To increase accessibility and broaden the potential user base, a Python interface to the core optimization routines of evortran is provided through a separate Fortran fpm package called pyevortran. This interface allows users to take advantage of the performance of Fortran-based GAs while working within Python environments.
The pyevortran package is a minimal Fortran project that uses evortran as a dependency. It contains only two modules: the module pyevortran__evolutions_float provides a C-binding wrapper for the evolve_population routine, and the module pyevortran__migrations_float provides a C-binding wrapper for the evolve_migration routine. So far only the routines acting on float populations are implemented. The two routines are compiled into a shared library which is then used by the Python package pyevortran that is installed using pip. Within this Python package, the shared library is accessed using Python’s ctypes module to expose the two functions evolve_population and evolve_migration. These functions can be called from Python to optimize arbitrary Python functions of the form f(x), where x is a NumPy array.
The Python interface currently supports only the two mentioned high-level optimization routines. Nevertheless, these routines allow the use of a variety of GA configurations by specifying optional arguments, just as in the native Fortran implementation. Another limitation to note is performance: when using the Python interface, the fitness function is typically written in Python and may run significantly slower than a compiled Fortran equivalent. If the evaluation of the fitness function dominates runtime of the GA, it may be worthwhile to translate the function into Fortran and use the native evortran package for optimal performance.
For installation, the pyevortran interface requires a Python v.3 installation, the gfortran Fortran compiler, and fpm version $\geq$ 0.12.0, as earlier versions do not support building shared libraries. The version of fpm currently distributed via PyPI is outdated (currently version 0.10.0) and will not work. To install a suitable version of fpm, one can download a recent binary from the GitHub releases page [60], make it executable, and move it to a directory in the system’s PATH, e.g.:
1chmod +x fpm
2mv fpm ~/.local/bin/
In addition, patchelf must be installed to embed runtime library paths (RPATH) into the shared libraries, ensuring that Python can find them at runtime without needing to manually set LD_LIBRARY_PATH. On Debian-based systems, patchelf can be installed with:
1sudo apt install patchelf
With the above mentioned prerequisites installed, one can proceed to install pyevortran. The first step is to clone the repository and navigate to its main folder:
1git clone https://gitlab.com/thomas.biekoetter/pyevortran
2cd pyevortran
Then one must source one of two provided environment setup scripts. As for evortran, see the discussion in Section 3.1.2, the script exports_debug.sh is intended for development and enables runtime checks and assertions, while exports.sh is optimized for production use by enabling compiler optimizations and disabling checks. These scripts also set up the number of OpenMP threads. To use the debug version, run:
1source exports.debug.sh
Alternatively, to use the optimized configuration, run:
1source exports.sh
Once the environment is configured, pyevortran can be built and installed with:
1make pyevortran
The makefile invokes fpm to build the shared library, uses pip to install the Python wrapper into the current environment, and finally applies patchelf to embed the necessary RPATHs into the shared libraries. After successful installation, users can import pyevortran in Python and use the evolve_population and evolve_migration functions to apply GAs to their own objective functions.
To test the installation, and to demonstrate the usage of pyevortran, a simple test program test_rastrigin.py is provided in the python/test folder:
1import numpy as np
2from pyevortran.evolutions import evolve_population
3
4A = 10
5def rastrigin(x):
6 return A * len(x) + np.sum(x**2 - A * np.cos(2 * np.pi * x))
7
8print()
9print("Minimizing Rastrigin function:")
10print()
11for dim in range(2, 11):
12 print(" Number of dimensions =", dim)
13 xmin = evolve_population(
14 rastrigin, dim,
15 pop_size=1000,
16 lower_lim=-5.12, upper_lim=5.12,
17 max_generations=1000,
18 fitness_target=1e-9, verbose=False,
19 selection="rank", selection_size=100,
20 elite_size=100)
21 print(" Minimum at xmin =", xmin)
22 print(" f(xmin) =", rastrigin(xmin))
23 print()
This script demonstrates the use of the evolve_population function from the Python interface to minimize the Rastrigin function, see Section 4.1.1 for details. The script performs minimization for dimensions ranging from two to ten and prints both the location of the minima and the corresponding function values. We note that only the first two arguments, the fitness function (rastrigin) and the dimensionality of the problem (dim), are required. All other arguments are optional and allow the user to specify GA parameters such as population size, selection strategy, elite size, and termination criteria.
@@ 4.1.2 Michalewicz function
1best_ind = evolve_population( &
@@ 4.1.3 Himmelblau function
$(x,y)$ (3.0, 2.0)$\approx$(3.0,\ 2.0),≈ ( 3.0 , 2.0 ) , (11)
$(x,y)$ (2.805, 3.131)$\approx$(-2.805,\ 3.131),≈ ( - 2.805 , 3.131 ) , (12)
$(x,y)$ (3.779,3.283)$\approx$(-3.779,\ -3.283),≈ ( - 3.779 , - 3.283 ) , (13)
$(x,y)$ (3.584,1.848)$\approx$(3.584,\ -1.848),≈ ( 3.584 , - 1.848 ) , (14)
@@ 4.1.4 Drop-Wave function
To demonstrate the flexibility of the evortran library, and to compare the performance of different crossover methods, we conducted a comprehensive set of runs using the evolve_migration function to minimze the Drop-Wave function using different crossover routines. For each dimensionality from $n=2$ to $n=10$, we tested each crossover operator (blend, sbx, one-point, two-point, uniform) in combination with all possible permutations of the three available selection routines (tournament, rank, roulette) and the three mutation strategies (uniform, shuffle, gaussian). Each unique combination was run 10 times, resulting in a total of 3310=903$\cdot$ 3$\cdot$ 10=903 ⋅ 3 ⋅ 10 = 90 runs per crossover method and dimension. The GA was configured via the following call:
@@ 4.2.1 Confronting extended Higgs sector with LHC data
    Perturbative unitarity: We verify that the leading-order $\to 2$ 2 scalar scattering amplitudes remain below the threshold $8\pi$ in the high-energy limit. This results in a set of inequalities on combinations of the scalar quartic couplings, whose precise form can be found in Ref. [16].
@@
Since the scan covers all physically distinct combinations of the three scalar mixing angles $\alpha_{1,2,3}$, see Table 4, no distinction between $h_{2}$ and $h_{3}$ is imposed on their possible masses and couplings. As a result, the orange points corresponding to the state $h_{2}$ and the green points corresponding to the state $h_{3}$ show a similar distribution in the two plots of Fig. 8. Furthermore, the mass ranges for $h_{2}$ and $h_{3}$ span from about 63 GeV to 1 TeV. The lower bound arises because masses below 62.562.562.5 GeV would open the decay channels $h_{1}\to h_{2}h_{2}/h_{3}h_{3}$, which tends to spoil the Standard-Model-like nature of $h_{1}$ unless the corresponding couplings governing these decays are fine-tuned to suppress the decays. The GA does not converge to such tuned regions of parameter space in this broad scan. On the upper end, the scan reaches the maximum mass range of 1 TeV set in the input.
@@
In the right plot of Fig. 9, we show the values of the coupling modifiers $\kappa_{u}$ vs. $\kappa_{d}$, which specifically correspond to the ratios of the $h_{1}$ couplings to top and bottom quarks, respectively, compared to the Standard Model. Unlike in the Yukawa Type I, these two couplings are now controlled independently via the alignment parameters $\xi_{u}$ and $\xi_{d}$. The distribution shows that most viable parameter points cluster in the region where both $\kappa_{u}$ and $\kappa_{d}$ are close to either +1+1+ 1 or 1-1- 1, reflecting Standard-Model-like behavior. However, one point clearly stands out as it corresponds to a parameter point where $\kappa_{u}$ and $\kappa_{d}$ have opposite signs. The two points in the left and the right plot of Fig. 9 corresponding to this parameter point are highlighted in orange. This sign configuration between $\kappa_{u}$ and $\kappa_{d}$ (also referred to as wrong-sign Yukawa coupling regime) is a distinctive feature of the flavor-aligned S2HDM compared to the S2HDM Type I and can, as observed here, still be consistent with current data, though it is associated with modified interference effects in loop-induced Higgs processes (such as gluon fusion production or the $h\to\gamma\gamma$ decay).
@@ 5 Conclusions
With its modular design and simple user interface, evortran enables users to easily apply evolutionary strategies to complex optimization problems, including those with non-differentiable, discontinuous, or noisy fitness functions. The library supports customization of GA components, native real and integer encodings, and parallel execution via OpenMP. evortran is installed with the Fortran package manager fpm, ensuring a straightforward dependency management, compilation and installation process, as well as a seamless integration into both simple scripts and larger code bases. To further enhance accessibility, evortran provides Python bindings available at:
https://gitlab.com/thomas.biekoetter/pyevortran.
This interface allows users to run the core optimization routines of evortran directly from Python.